# Event-based backpropagation on the neuromorphic platform SpiNNaker2

### v1

@@ 1 Introduction
Gradient-based learning using neuromorphic hardware has been previously implemented using “in-the-loop” approaches, where neuromorphic hardware computes spike times and state variables in an inference phase (the forward pass) while a classical computer computes gradients and implements backpropagation for the computation of surrogate gradients [Cramer et al., 2022, Esser et al., 2015] or exact gradients [Göltz et al., 2021, Pehle et al., 2023]. Previous work implemented the E-prop algorithm [Bellec et al., 2020] on SpiNNaker1 [Perrett et al., 2022] and an FPGA-based prototype of SpiNNaker2 [Rostami et al., 2022]. This algorithm computes surrogate gradients using local eligibility traces and a global error signal, without error backpropagation through network layers, which might alter scalability. Besides SpiNNaker, the GeNN code generation framework [Knight et al., 2021] has been used to create a GPU-based implementation of EventProp [Nowotny et al., 2022].
@@ 2.2 Programs on Processing Elements
A common regularization method in machine learning is to process a subset of the training batch in parallel: gradients are averaged across a mini-batch [Goodfellow et al., 2016]. Mini-batch training can achieve better generalization while speeding up training by processing training examples in a given mini-batch in parallel. In the case of training on Spinnaker2, we can take advantage of the inherently parallel nature of the architecture to deploy identical copies of the network on multiple cores. Different copies of a network process different input samples, implementing mini-batch training. Except for the optimisation program, all described programs are duplicated on the chip for each training example processed in parallel.
@@
Neuron Layer Program: This program simulates a layer of leaky integrate-and-fire neurons by computing the forward and backward pass. It uses a clock-driven simulation based on the discretized differential equations and their adjoints [Wunderlich and Pehle, 2021]. In a given time step, different network layers are computed asynchronously on different PEs and event packets representing spikes (forward pass) or error signals (backward pass) are transmitted by the network-on-chip and buffered by the receiving PEs to update variables in the next time step. PEs store dense weight matrices (in contrast to the original synapse-based implementation of SpiNNaker) and use incoming spikes’ payloads to retrieve the synaptic weight corresponding to the arriving spike and to update the synaptic current of the target neuron.
During the backward pass, spikes are distributed in reverse order, and carry state variables representing error signals within the payload of event packets. At this point, the network runs in “reverse-time” and follows the adjoint dynamics, with error signals being transmitted at the spike times computed during the forward pass. In this way, the backward pass implements a spiking neural network with graded spikes (representing the error signal) at fixed times, making it suited for an event-based implementation on the neuromorphic platform. The backward pass implements a discretization of the adjoint system [Wunderlich and Pehle, 2021], following an “optimize-then-discretize” approach. In this way, the program uses event-based backpropagation to compute a discretized approximation to the exact gradients of the simulated spiking neural network. The complete set of discretized equations is provided in section 6.1.
@@ 3 Results
In addition to batch-parallelised processing, we report on results where the networks are trained on a single sample at a time, for a single epoch, on a limited training-set of samples (Figure 3, right). Such a scenario is particularly relevant for on-line learning on neuromorphic hardware, where the network is trained on a continuous stream of data, which would be critical for embedded systems and autonomous agents. Again, the results show a close match between on-chip and off-chip simulations, and satisfying accuracy considering the heavy limitations.
@@ 5 Acknowledgement
This project is partially funded by the EIC Transition under the ”SpiNNode” project (grant number 101112987), and by the Federal Ministry of Education and Research of Germany in the programme of “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KISK001K. This work was partially funded by the German Federal Ministry of Education and Reseach (BMBF) within the KI-ASIC project (16ES0996), and by the BMBF and the free state of Saxony within the ScaDS.AI center of excellence for AI research. The authors would like to acknowledge the 2022 Telluride Neuromorphic Cognition Engineering Workshop where this project was initiated.

### v4

@@ 1 Introduction
Gradient-based learning using neuromorphic hardware has been previously implemented using “in-the-loop” approaches, where neuromorphic hardware computes spike times and state variables in an inference phase (the forward pass) while a classical computer computes gradients and implements backpropagation for the computation of surrogate gradients [Cramer et al., 2022, Esser et al., 2015] or exact gradients [Göltz et al., 2021, Pehle et al., 2023]. While we were completing this project, another approach to on-chip backpropagation has also been proposed, implemented on Intel’s Loihi chip [Renner et al., 2024]. Previous work implemented the E-prop algorithm [Bellec et al., 2020] on SpiNNaker1 [Perrett et al., 2022] and an FPGA-based prototype of SpiNNaker2 [Rostami et al., 2022]. This algorithm computes surrogate gradients using local eligibility traces and a global error signal, without error backpropagation through network layers, which might alter scalability. Besides SpiNNaker, the GeNN code generation framework [Knight et al., 2021] has been used to create a GPU-based implementation of EventProp [Nowotny et al., 2022].
@@ 2.2 Programs on Processing Elements
A common parallelization method in machine learning is to process a subset of the training batch in parallel: gradients are averaged across a mini-batch [Goodfellow et al., 2016]. Mini-batch training can achieve better generalization while speeding up training by processing training examples in a given mini-batch in parallel. In the case of training on SpiNNaker2, we can take advantage of the inherently parallel nature of the architecture to deploy identical copies of the network on multiple cores. Different copies of a network process different input samples, implementing mini-batch training. Except for the optimisation program, all described programs are duplicated on the chip for each training example processed in parallel.
@@
Neuron Layer Program: This program simulates a layer of leaky integrate-and-fire neurons by computing the forward and backward pass. It uses a clock-driven simulation based on the discretized differential equations and their adjoints [Wunderlich and Pehle, 2021]. In a given time step, different network layers are computed asynchronously on different PEs and event packets representing spikes (forward pass) or error signals (backward pass) are transmitted by the network-on-chip and buffered by the receiving PEs to update variables in the next time step. PEs store dense weight matrices (in contrast to the original synapse-based implementation of SpiNNaker2) and use incoming spikes’ payloads to retrieve the synaptic weight corresponding to the arriving spike and to update the synaptic current of the target neuron.
During the backward pass, error signals are propagated in reverse order through event-driven floating point messages within the payload of event packets. While the underlying mechanism shares the temporal event-driven nature of spike-based computation, these backward messages differ from biological spikes as they carry signed floating point values rather than binary spike events. At this point, the network runs in “reverse-time” and follows the adjoint dynamics, with error signals being transmitted at the same time points computed during the forward pass. This implementation leverages the event-based architecture of the neuromorphic platform while acknowledging that the backward pass operates on continuous-valued error gradients rather than discrete spikes. The backward pass implements a discretization of the adjoint system [Wunderlich and Pehle, 2021], following an “optimize-then-discretize” approach. In this way, the program uses event-driven backpropagation to compute a discretized approximation to the exact gradients of the simulated spiking neural network. The complete set of discretized equations is provided in section 6.1.
@@ 3 Results
In addition to batch-parallelised processing, we report on results where the networks are trained on a single sample at a time, for a single epoch, on a limited training-set of samples (Figure 3, right). While our current implementation processes the entire 30ms input sequence before performing the backward pass (total 61ms per sample), the neuromorphic architecture offers opportunities for optimization through pipelining. Specifically, the backward pass for sample could be computed in parallel while processing the forward pass of sample , effectively reducing the per-sample processing time closer to 30ms. This pipelined approach would maintain the exact gradient computation while significantly reducing the effective latency between receiving an input and applying the corresponding weight updates. The method still requires maintaining the full trajectory of the forward pass to compute precise updates, but the parallel processing would enhance its applicability for systems requiring rapid adaptation. Such a scenario is particularly relevant for online learning on neuromorphic hardware, where the network is trained on a continuous stream of data, which would be critical for embedded systems and autonomous agents. The results show a close match between on-chip and off-chip simulations with satisfying accuracy considering the heavy limitation of this training regime.
@@ 5 Acknowledgement
This project is partially funded by the EIC Transition under the ”SpiNNode” project (grant number 101112987), and by the Federal Ministry of Education and Reseach of Germany in the programme of ”Souverän. Digital. Vernetzt.” Joint project 6G-life, project identification number: 16KISK001K. This work was partially funded by the German Federal Ministry of Education and Reseach (BMBF) within the KI-ASIC project (16ES0996), and by the BMBF and the free state of Saxony within the ScaDS.AI center of excellence for AI research. The authors would like to acknowledge the 2022 Telluride Neuromorphic Cognition Engineering Workshop where this project was initiated.